# Quantum latent distributions in deep generative models

### v1

@@ Abstract
Many successful families of generative models leverage a low-dimensional latent distribution that is mapped to a data distribution. Though simple latent distributions are commonly used, it has been shown that more sophisticated distributions can improve performance. For instance, recent work has explored using the distributions produced by quantum processors and found empirical improvements. However, when latent space distributions produced by quantum processors can be expected to improve performance, and whether these improvements are reproducible, are open questions that we investigate in this work. We prove that, under certain conditions, these "quantum latent distributions" enable generative models to produce data distributions that classical latent distributions cannot efficiently produce. We also provide actionable intuitions to identify when such quantum advantages may arise in real-world settings. We perform benchmarking experiments on both a synthetic quantum dataset and the QM9 molecular dataset, using both simulated and real photonic quantum processors. Our results demonstrate that quantum latent distributions can lead to improved generative performance in GANs compared to a range of classical baselines. We also explore diffusion and flow matching models, identifying architectures compatible with quantum latent distributions. This work confirms that near-term quantum processors can expand the capabilities of deep generative models.
@@ I Introduction
In recent years, several successful families of deep generative models have emerged that learn to map samples from a latent distribution, a typically low-dimensional probability distribution, to a higher-dimensional data distribution. For instance, models that use latent distributions such as generative adversarial networks (GANs) [1, 2], latent diffusion models [3, 4], flow models [5, 6] and variational autoencoders (VAEs) [7, 8] have all achieved remarkable performance on a wide range of datasets from images to proteins. During training, these models learn a mapping between the structure of the latent distribution and that of the data. As such, the structure of this latent distribution plays a critical role. For instance, in GANs, making the structure of the latent distribution used by the generator correspond to that of the data has been shown to significantly improve performance [9, 2, 10].
The design space in which these new algorithms have been developed has typically been limited to the range of functions that are efficient to implement on modern CPUs and GPUs. As such, relatively simple latent distributions are typically used, for example a multivariate Gaussian distribution transformed by a small neural network [2]. However, the use of simple latent distributions with finite-capacity neural network architectures can be a limiting factor for modeling complex datasets [10]. As an example, many quantum processes cannot be efficiently simulated using classical (i.e. non-quantum) computational methods [11]. Learning the data distribution arising from such a process is thus likely to be a challenging task for classical generative models using a simple latent distribution.
@@
In this work, we develop a deeper understanding of the use of quantum distributions in the latent space, referred to as "quantum latent distributions", and perform extensive experiments to benchmark this approach. We theoretically investigate the relationship between the complexity class of the latent distribution and that of the generated distribution. We prove that with a quantum latent distribution some types of neural network architectures can produce a wider range of distributions in the data space than could be achieved with classical latent distributions. This can allow a model to achieve improved performance on some datasets. We develop an intuition for why and when this improved performance can be achieved in practice. Based on this theory, we select an experimental setting in which we benchmark a range of latent distributions and GAN models. We find that a quantum latent distribution can indeed lead to improved generator performance compared to a relevant range of classical distributions. We then propose avenues for extending the use of a quantum latent distributions beyond GANs to diffusion and flow matching models.
@@
  • We show that, under some assumptions on the generative model, using a quantum distribution in the latent space can broaden the range of output distributions and lead to improved performance on some datasets.
  • We benchmark this approach on both a toy quantum dataset and on the QM9 quantum chemistry dataset, and show that a quantum latent distribution in a GAN can improve model performance compared to a range of relevant classical distributions.
  • We identify families of deep generative models beyond GANs that are compatible with the use of a quantum latent distribution.
While the benefits of using a quantum latent distribution are dataset and model dependent, our results validate the notion that these distributions can be a useful tool to improve the performance of deep generative models.
@@ II Background and Related Work
The impact of the latent distribution on the performance of a generative model has been a significant topic of investigation. The choice of latent distribution has been empirically found to have a significant impact on model performance in different types of architectures, including GANs [2, 21, 22, 9, 23, 24, 25], VAEs [26, 8], and flow matching models [27]. Some general arguments have been put forward to explain why, for example, uncorrelated latent distributions are not well suited for capturing correlated features in complex data [2]. However, developing a more general theoretical understanding of the relationship between the latent distribution and the data distribution has been challenging. To address this issue, [10] introduced a novel distance between these two spaces, which is minimized when the latent most effectively capitalizes on the capacity of the generator neural network. They harness this insight to increase model performance using a two-stage decoupled autoencoder (DAE) architecture that shapes the latent distribution. In our work, we extend this formalism to develop an understanding of the impact of the computational complexity of the latent distribution on model performance.
There have been several previous investigations of quantum latent distributions in generative models with a focus on GANs, in which empirical performance improvements were observed. However, much of this work focused on demonstrating feasibility, with a smaller emphasis on benchmarking or on developing theoretical foundations for this approach. In [28] a quantum latent distribution is used within a generative algorithm to propose new molecules for cancer therapy, and in [19] this approach is used for image reconstruction. In [17, 18] and [16, 12], small-scale quantum latent distributions in GANs are found to outperform a classical latent distribution on the QM9 and MNIST datasets. However, in those works the use of a trained quantum distribution compared to an untrained classical baseline makes it difficult to draw generalizable conclusions. [20] empirically show an improvement in the BigGAN model [23] using a range of untrained quantum latent distributions, but do not benchmark against other classical latent distributions or provide theoretical foundations for this approach. We go beyond this prior work both by developing a theoretical understanding of the use of quantum distributions, and by benchmarking against a range of classical distributions in an apples-to-apples experimental setting.
@@ II.1 Quantum circuit sampling
Some classes of discrete distribution sampling tasks can be performed faster on near-term quantum computers than on classical computers [15]. Two specific realizations of quantum computers that can perform such tasks are boson sampling [11, 29] and random circuit sampling [30, 13]. In boson sampling, identical single photons are sent into a random interference circuit, and measurements of the output locations of these photons provide samples from a classically-intractable probability distribution. Similarly, in random circuit sampling, a random sequence of gates is applied to qubits, and the output measurements of the qubit states are also a provably hard-to-sample distribution.
The experiments we perform in this work use photonic boson sampling distributions [11]. Photonic quantum systems have several advantages compared to other quantum computing modalities that motivate using them to generate latent distributions. Though they are not universal for quantum computation, they naturally produce highly non-uniform distributions [31] that differ significantly from commonly used classical latent distributions. Moreover, photonic systems are not subject to decoherence, which pushes the outputs of quantum circuits implemented with other modalities closer to a classical uniform distribution [32, 20]. The advantages of photonics were leveraged in [33, 34] to perform some the largest-scale demonstrations of quantum machine learning to date.
@@ III.1 Preliminaries
Sampling problems: We consider the problem of drawing samples from a given probability distribution. We define $\mathcal{C}$ as the class of probability distributions that can be approximately sampled in Poly($n,1/\epsilon$) time using classical computers. Here, $n$ represents the dimension of the samples to be generated, $\epsilon$ the error within which we sample from this distribution. Quantum computers can sample from some distributions that cannot be efficiently sampled from classically [15]. We thus define $\mathcal{Q}$ as the class of distributions that can be sampled in Poly($n,1/\epsilon$) with a quantum computer, but not with a classical computer. We provide rigorous definitions of these classes in the appendix.
In this section, we consider a family of neural network architectures $G$ mapping from $R^{d_{z}}$ to $R^{d}$, where $d_{z}$ is the latent space dimension and $d$ is the data space dimension with $d_{z}\leq d$. We assume that these architectures have bounded complexity, where complexity is loosely defined in terms of width or depth of the network, such that the output can be efficiently produced from the input.
@@ III.2 Computational complexity of the latent distribution
Our first contribution is making a connection between the complexity class of the latent distribution and the complexity class of the output distribution of a deep generative model. We start with a simple observation about the complexity of pushforward distributions.
@@
Conversely, if a latent distribution $P_{z}\in\mathcal{Q}$, then the pushforward distribution is not necessarily in $\mathcal{Q}$. There are scenarios in which an initial quantum distribution is converted to a classical distribution, for example if the neural network is trained to produce constant outputs. A relevant question is whether this collapse from $\mathcal{Q}$ to $\mathcal{C}$ is expected to be a common occurrence or the exception.
Here, we derive sufficient conditions under which $P_{g(z)}$ cannot be efficiently sampled from classically:
@@
Though these assumptions may appear restrictive, they apply to some types of deep neural network architectures. As an example, consider a deep feedforward neural network mapping a small-dimensional latent space to a larger-dimensional data space through layers of increasing width with a LeakyReLU activation. Since by assumption no linear layer decreases the size of the inputs, the input of each linear layer can in general be efficiently reconstructed from the outputs by inverting a set of linear equations. The use of LeakyReLU makes the entire network invertible. Moreover, feedforward neural networks are typically Lipschitz-continuous [35]. This architecture thus satisfies the assumptions. In our benchmarking on the QM9 dataset, we adopt a model architecture that closely matches this example.
Beyond this simple example, we note that other types of neural networks such as convolutional neural networks are also generally Lipschitz continuous [35]. A notable exception is dot product self-attention, though alternatives that are Lipschitz continuous have been developed [36]. Moreover, the Lipschitz constant of a neural network is sometimes used as a proxy for model complexity [37]. Since we consider models with bounded complexity, the Lipschitz continuity assumption is thus often satisfied in practice for a wide range of architectures.
Even when a neural network does not strictly satisfy the conditions above, the output distribution may still be non-classical for a range of possible input quantum distributions. Theorem 1 provides a theoretical basis for suggesting that many modern neural network architectures can in practice use latent distributions within $\mathcal{Q}$ to produce a broader set of distributions than would be possible only with distributions within $\mathcal{C}$. This is illustrated in figure 1.
@@ III.3 Impact on model performance
In this section, we discuss the implications to model performance, defined in terms of GAN-induced distance, through the use of a quantum latent distribution. First, we show that at least some data distributions can be better approximated using a quantum latent distribution:
@@
This shows that the class of pushforward distributions achieved by a quantum latent distribution under certain conditions cannot be approximated with any choice of classical latent distribution. Intuitively, the corollary demonstrates that there exist classes of generative tasks for which the performance we can achieve with classical latent distributions will be strictly worse than that which can be achieved with quantum latent distributions. This may include datasets that result from a classical transformation of an underlying quantum process.
We can now reason in terms of performance on different types of datasets. For a target data distribution that is in $\mathcal{C}$, we can distinguish between two cases illustrated in figure 1. First, if this distribution is in the pushforward distribution of classical latents that are also in $\mathcal{C}$, then there is no theoretical advantage to using a quantum latent. In the second case, the target distribution is not in the pushforward distribution of classical latents. Since we consider generators with finite capacity and latent spaces that are smaller than the data space, this may be a common scenario. In that case, some achievable quantum pushforward distributions will be closer than any achievable classical pushforward distributions in terms of GAN-induced distance. As such, a quantum latent distribution may lead to an advantage even for a classical data distribution.
@@ III.4 Achieving an advantage in practice
The results above establish a theoretical basis for the use of quantum distributions in generative models, but the conditions described above for achieving a performance advantage over classical distributions are still highly abstract. Here, we propose two mechanisms through which quantum distributions can be helpful, which can provide practical guidance as to when to use such a distribution.
First, the use of a quantum latent distribution can endow a model with a beneficial "inductive bias" [38] for certain types of dataset. A quantum latent distribution can be expected to provide such a bias for a dataset arising from a fundamental physical process that is more efficiently described by quantum mechanics than by classical mechanics. Moreover, as suggested by corollary 1 the performance of models with classical latent distributions is inherently limited for such datasets. This leaves a broader scope for a performance improvement with a quantum distribution.
Figure 2: Comparison between GANs trained on a 2D mixture of Gaussians dataset with different latent distributions. The four latent distributions represent a progression from the commonly used Gaussian distribution (left) to a quantum distribution (boson sampler, right) with more information provided in section 4. The main failure mode is a tendency to interpolate between different modes, and the model with the quantum latent distribution is least affected. This could be explained by the fact that quantum distributions cannot be tractably factorized into independent factors of variation that could encourage smooth interpolations.
Second, a quantum latent distribution prevents a model from learning a fully factorized representation of the data, which we hypothesize can be helpful for multimodal datasets. A quantum distribution that is classically hard to simulate cannot be efficiently factorized into independent factors of variation. Though factorizable representations are often viewed positively since they are more interpretable, empirical evidence for other benefits are mixed [39, 40]. Factorized representations can perform poorly on multimodal datasets without additional modifications [41]. Though factorized representations have been shown to facilitate smooth interpolation in the latent space [42], when the data is multimodal a representation that encourages interpolation between the modes may be detrimental. To further investigate, we trained GANs with different latent distributions on a simple multimodal 2D mixture-of-Gaussians dataset. Our results are shown in figure 2 and support this hypothesis, where we observe that the quantum distribution produces the model that interpolates the least between the modes. Further empirical evidence for improved data coverage from quantum latent distributions was also observed in [20].
Taken together, these mechanisms provide guidelines for when quantum latent distributions may provide a practical advantage. Though we do not expect them to provide a benefit for all models and datasets (with negative examples provided in appendix H), quantum latent distributions may be well suited for datasets where the underlying data generating process is poorly described by classical mechanics, or datasets that are highly multimodal. To empirically verify this intuition, in the next section we perform experiments focusing on datasets arising from physics and chemistry, where we compare the performance of different latent distributions on several metrics including data diversity.

### v2

@@ Abstract
Many successful families of generative models leverage a low-dimensional latent distribution that is mapped to a data distribution. Though simple latent distributions are often used, the choice of distribution has a strong impact on model performance. Recent experiments have suggested that the probability distributions produced by quantum processors, which are typically highly correlated and classically intractable, can lead to improved performance on some datasets. However, when and why latent distributions produced by quantum processors can improve performance, and whether these improvements are connected to quantum properties of these distributions, are open questions that we investigate in this work. We show in theory that, under certain conditions, these "quantum latent distributions" enable generative models to produce data distributions that classical latent distributions cannot efficiently produce. We provide intuition as to the underlying mechanisms that could explain a performance advantage on real datasets. Based on this, we perform extensive benchmarking on a synthetic quantum dataset and the QM9 molecular dataset, using both simulated and real photonic quantum processors. We find that the statistics arising from quantum interference lead to improved generative performance compared to classical baselines, suggesting that quantum processors can play a role in expanding the capabilities of deep generative models.
@@ I Introduction
In recent years, several successful families of deep generative models have emerged that learn to map samples from a latent distribution, a typically low-dimensional probability distribution, to a higher-dimensional data distribution. For instance, models that use latent distributions such as generative adversarial networks (GANs) [21, 28], latent diffusion models [50, 9], and flow matching models [36, 4] have all achieved remarkable performance on a wide range of datasets from images to proteins. During training, these models learn a mapping between the structure of the latent distribution and that of the data. As such, the structure of this latent distribution plays a critical role. For instance, in both GANs [5, 28, 25] and flow matching models [34, 32], matching the structure of the latent distribution used by the generator to that of the data has been shown to significantly improve performance.
The design space for these algorithms has typically been limited to the range of functions that are efficient to implement on modern CPUs and GPUs. As such, relatively simple latent distributions are typically used, for example a multivariate Gaussian distribution transformed by a neural network [28]. However, the use of simple latent distributions with finite-capacity neural network architectures can be a limiting factor for modeling complex datasets [25]. As an example, many quantum processes cannot be efficiently simulated using classical (i.e. non-quantum) computational methods [1]. Learning the data distribution arising from such a process is thus likely to be challenging for classical generative models using a simple latent distribution.
@@
In this work, we develop a deeper understanding of the use of quantum latent distributions, and perform experiments to benchmark this approach. We theoretically investigate the relationship between the complexity class of the latent distribution and that of the generated distribution. We show that, under some assumptions on the neural network architecture, a quantum latent distribution can produce a wider range of distributions in the data space than could be achieved with classical latent distributions. This can allow a model to achieve improved performance on some datasets. We develop an intuition for why and when this improved performance can be achieved in practice. Based on this intuition, we select a synthetic quantum dataset and the QM9 chemistry dataset for benchmarking experiments. We compare a quantum distribution produced by quantum interference between photons to a range of classical distributions, including a photonic distribution where we switched off quantum interference effects. We find that the quantum distribution outperforms the other distributions on these datasets, indicating that some properties of quantum distributions can indeed lead to improved generative performance.
@@
  • We show that, under some assumptions on the generative model, using a quantum distribution in the latent space allows the generative model to produce distributions that could not be efficiently produced classically.
  • We provide intuition as to when and why a quantum distribution can lead to improved performance on a real dataset.
  • We benchmark this approach on both a synthetic quantum dataset and the QM9 quantum chemistry dataset, and show that a quantum latent distribution in a GAN can improve model performance compared to a range of relevant classical distributions.
While performance advantages yielded by a quantum latent distribution are dataset and model dependent, our results validate the notion that these distributions can be a useful tool to improve the performance of deep generative models.
@@ II Background and Related Work
Figure 1: Our work provides a theoretical and empirical study of the use of non-classical distributions produced by quantum processors in the latent space of generative models. Even on a simple dataset such as the 2D mixture of Gaussians shown here, training a simple GAN with different latent distributions yields very different results. We compare four latent distributions representing a progression from the commonly used Gaussian distribution (bottom) to a quantum distribution produced by a photonic quantum processor (top). The main failure mode is a tendency to interpolate between different modes in the data, and the model with the quantum latent distribution is least affected. Further information on this experimental setting can be found in Appendix C. In this paper, we investigate this approach from a theoretical perspective, propose some intuition to explain these results, and conduct benchmarking experiments to compare quantum distributions to relevant classical distributions in an apples-to-apples setting.
The impact of the latent distribution on the performance of a generative model has been a significant topic of investigation. The choice of latent distribution has a significant impact on model performance in different types of architectures, including GANs [28, 8, 44, 5, 10, 62, 38] and flow matching models [55, 32]. Some general arguments have been put forward to explain why, for example, uncorrelated latent distributions are not well suited for capturing correlated features in complex data [28]. However, developing a more general theoretical understanding of the relationship between the latent distribution and the data distribution has been challenging. To address this issue, [25] introduced a novel distance between these two spaces, which is minimized when the latent most effectively capitalizes on the capacity of the generator neural network. In our work, we extend this formalism to develop an understanding of the impact of the computational complexity of the latent distribution on model performance.
There have been several previous investigations of quantum latent distributions in generative models with a focus on GANs, in which empirical performance improvements were observed. However, much of this work focused on demonstrating feasibility, with a smaller emphasis on benchmarking or on developing theoretical foundations for this approach. In [57] a quantum latent distribution is used within a generative algorithm to propose new molecules for potential medical applications, and in [63] this approach is used for image reconstruction. In [35, 27] and [61, 51], small-scale quantum latent distributions in GANs are found to outperform a classical latent distribution on the QM9 and MNIST datasets. However, in these works the use of a trained quantum distribution compared to an untrained classical baseline makes it difficult to draw generalizable conclusions. [26] empirically show an improvement in the BigGAN model [10] using a range of untrained quantum latent distributions, but do not benchmark against other classical latent distributions or provide theoretical foundations for this approach. We go beyond this prior work both by developing a theoretical understanding of the use of quantum distributions in these models, and by benchmarking against a range of classical distributions in an apples-to-apples experimental setting.
@@ II.1 Quantum circuit sampling
Some classes of discrete distribution sampling tasks can be performed faster on near-term quantum computers than on classical computers [24]. Two specific realizations of quantum computers that can perform such tasks are random circuit sampling [43, 6] and boson sampling [1, 66]. Our work uses photonic boson sampling systems, in which identical single photons are sent into a random interference circuit, creating an entangled state. Measurements of the output locations of these photons provide samples from a classically-intractable probability distribution.
The use of boson sampling distributions to investigate the impact of a quantum distribution in the latent space is motivated by several considerations. First, though they are not universal for quantum computation, they naturally produce highly non-uniform distributions [2] that differ significantly from commonly used classical latent distributions. Moreover, photonic systems are subject to photon loss but not to decoherence. In other platforms, decoherence pushes the outputs of quantum circuits towards a classical uniform distribution [60], making it harder to observe a difference in performance [26]. These properties of photonic systems were leveraged in [12, 65] to perform some the largest-scale demonstrations of quantum machine learning to date.
@@ III.1 Preliminaries
Sampling problems: We consider the problem of drawing samples from a given probability distribution. We define $\mathcal{C}$ as the class of probability distributions that can be approximately sampled in Poly($n,1/\epsilon$) time using classical computers. Here, $n$ represents the dimension of the samples to be generated, and $\epsilon$ the error within which we sample from this distribution. We also define $\mathcal{Q}$ as the class of distributions that can be sampled in Poly($n,1/\epsilon$) with a quantum computer, but not with a classical computer. We provide rigorous definitions of these classes in the appendix.
In this section, we consider a family of neural network architectures $G$ mapping from $\mathbb{R}^{d_{z}}$ to $\mathbb{R}^{d}$, where $d_{z}$ is the latent space dimension and $d$ is the data space dimension with $d_{z}\leq d$. Similarly to [25], we assume that these architectures have bounded complexity according to some measure, such that the output can be efficiently produced from the input. Possible metrics for complexity can include width, depth, or Lipschitz constant of the network.
@@ III.2 Computational complexity of the latent distribution
Our first contribution is to formally establish the conditions under which a quantum latent distribution remains non-classical after being transformed by a neural network. We start with a simple observation about the complexity of classical pushforward distributions.
@@
However, the pushforward distribution of a quantum distribution in $\mathcal{Q}$ is not necessarily in $\mathcal{Q}$. There are scenarios in which an initial quantum distribution is converted to a classical distribution, for example if the neural network produces constant outputs. Here, we establish sufficient conditions under which the pushforward distribution of a quantum distribution cannot be efficiently sampled from classically:
@@
We can explicitly construct a deep neural network architecture that provably satisfies these assumptions. Consider a multi-layer perceptron mapping a small-dimensional latent space to a larger-dimensional data space through linear layers of increasing width with an invertible LeakyReLU activation. For such a model, the inputs can in general be efficiently reconstructed from the inputs by inverting a set of linear equations. This architecture satisfies the assumptions. In our benchmarking on the QM9 dataset, we adopt a model architecture that matches this example.
Even when a neural network does not strictly satisfy these sufficient conditions, Theorem 1 provides a basis for suggesting that many modern neural network architectures can map a non-classical distribution to another non-classical distribution. For example, though most neural networks are not analytically invertible, the “generator inversion” problem of finding a latent vector that produces a given output can be tractable in practice [14, 3]. Moreover, a wide range of neural network architectures are Lipschitz-continuous [59, 11].
@@ III.3 Impact on model performance
Here, we discuss the implications to model performance, defined in terms of GAN-induced distance, through the use of a quantum latent distribution. First, we show that at least some data distributions can be better approximated using a quantum latent distribution:
@@
This shows that the class of pushforward distributions achieved by a quantum latent distribution under certain conditions cannot be approximated with any choice of classical latent distribution.
We can now reason in terms of performance on different types of datasets. For a target data distribution that is in $\mathcal{C}$, we can distinguish between two cases illustrated in figure 2. First, if this distribution is in the pushforward distribution of classical latents that are also in $\mathcal{C}$, then there is no theoretical advantage to using a quantum latent distribution. In the second case, the target distribution is not in the pushforward distribution of classical latents. Since we consider generators with finite capacity and latent spaces that are smaller than the data space, this may be a common scenario. In that case, some achievable quantum pushforward distributions will be closer than any achievable classical pushforward distributions in terms of GAN-induced distance. As such, a quantum latent distribution may lead to an advantage even for a classical data distribution.
@@ III.4 Achieving an advantage in practice
The results above establish a theoretical basis for the use of quantum distributions in generative models, but the conditions described above for achieving a performance advantage over classical distributions are still abstract. Here, we propose two mechanisms through which quantum distributions can lead to improved performance in practice.
First, a consequence of multi-particle entanglement is that quantum distributions cannot be factorized into independent factors of variation. This prevents a model from learning a factorized representation of the data. Though factorizable representations are often viewed positively since they are more interpretable, empirical evidence for other benefits are mixed [37, 41]. Factorized representations can perform poorly on multimodal datasets without additional modifications [56], and non-factorizable determinantal point processes have been integrated into GANs [19] and flow matching [42] to improve output diversity. We hypothesize that using a non-factorizable distribution can in some cases improve learning by biasing the training process towards more complex models for the data, avoiding oversimplified models that focus on uncorrelated features. Quantum distributions exhibit a strong form of non-factorizability: unlike classically tractable non-factorized distributions, which can still be transformed into factorized distributions when their probabilities are efficiently computable with methods such as cumulative transforms or rejection sampling, no such efficient transformation is possible in general for quantum distributions.
Second, quantum distributions are highly non-uniform, with strong correlations at multiple orders [2, 47]. We anticipate that this can provide a beneficial inductive bias for datasets that have similar properties. For example, a quantum latent distribution can be expected to provide such a bias for a dataset arising from a fundamental physical process that is better described by quantum mechanics than by classical mechanics. However, though non-uniformity and strong correlations are features of quantum distributions that may provide a benefit, we note that these are not uniquely quantum features and other classical distributions share similar properties. To help assess whether the specific statistics arising from quantum mechanics can be beneficial, our experiments in the next section compare the performance of a photonic quantum distribution to that of a photonic distribution where quantum interference effects are switched off.
To test these intuitions, we trained GANs with two hidden layers and different latent distributions (described in the next section) on a 2D mixture of Gaussians dataset. Though simple, this dataset is highly non-uniform and multimodal, making it a priori a good match for the properties of quantum latent distributions. Our results are shown in figure 1 and support these intuitions, where we observe that the quantum distribution produces the model that most closely matches the data. This is in line with prior work from [38], where up to 7 layers were required to successfully map a Gaussian latent distribution to a mixture of Gaussians.